A 2600 kg truck traveling to the north is slowed down uniformly from an initial velocity of 29 m/s by a 6350 N braking force acting opposite the truck's motion.

The truck's velocity after 1.6 seconds is   m/s.

The time is takes the truck to come to a complete stop is 


1
Expert's answer
2021-03-02T18:05:35-0500

(a)

"F\Delta t=m(v_f-v_i),""F\Delta t=mv_f-mv_i,""v_f=\dfrac{F\Delta t+mv_i}{m},""v_f=\dfrac{-6350\ N\cdot1.6\ s+2600\ kg\cdot29\ \dfrac{m}{s}}{2600\ kg}=25.1\ \dfrac{m}{s}."

(b)

"F\Delta t=m(v_f-v_i)""\Delta t=\dfrac{2600\ kg\cdot(0-29\ \dfrac{m}{s})}{-6350\ N}=11.9\ s."